find z if 3z-19=173-wang

Found 3 solutions by maribeth_abella, MathGrace, uksraj:
Answer by maribeth_abella(11)

3z - 19 = 173
3z = 173 + 19
3z = 192
z = 192/3
z = 64

Answer by MathGrace(6)

To solve any equation, you can use PEMDAS (Order of Operations) backwards. First get rid of any subtraction and addition.
Add 19 to both sides. 3z-19+19=173+19 --> 3z=192
Then get rid of any multiplication and division.
Divide both sides by 3. 3z/3=192/3 --> z=64
Finally, plug your answer back in to make sure you did it right.
3(64)-19=192-19=173 We have the right answer! :-)

Answer by uksraj(9)

3z-19=173
adding both sides 19 to get rid of 19
3z-19+19=173+19
3z=192
z=192/3
=62